blank + blank + blank = 21 lb.
How do I solve?

Found 2 solutions by jim_thompson5910, solver91311:
Answer by jim_thompson5910(35256)

x+x+x = 21

3x = 21

x = 21/3

x = 7

So

7+7+7 = 21

Answer by solver91311(24713)



Depends. Do all three of your blanks have the same value or is there some other relationship?

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism